If I adding odd numbers from -7 to 101. Can you show me how to get the sum the quickest way possible?
Answer by stanbon(75887)

If I adding odd numbers from -7 to 101. Can you show me how to get the sum the quickest way possible?
---
-7,-5,-3.........95,97,99
===============================
Add the 1st and the last to get 92
Add the 2nd and the next to last to get 92
-----
Add the 3rd and the 2nd to last to get 92
-----
The sum of each of the pairs is 92
---
How many pairs are there?
Find the number of terms using a(n)= a(1) + (n-1)d
99 = -7+(n-1)2
106 = (n-1)2
53 = n-1
n = 54
So there are 54 terms in the series.
So there are 27 pairs of terms.
-----
So the sum of all the pairs is:
27*92 = 2484
===================
A German mathematician named Euler figured
this out in grade school.
Formula: S(n) = (n/2)(first + last)
Try it with shorter series like 1,2,3,4
===================================
Cheers,
Stan H.